OpenStudy (princessaurora):

*Will fan and Medal* An experiment is set up to study the weights of 35 mice after they are injected with a drug. The population mean is 23.50 grams with a standard deviation of 3.40 grams. The standard error of the sample mean is BLANK . (Round off your answer to the nearest hundredth.) What is the blank?

OpenStudy (johnweldon1993):

Looks like the standard error of the sample mean is \(\large SE = \frac{\sigma}{\sqrt{n}}\) So here we have \(\large SE = \frac{3.40}{\sqrt{35}}\) Can someone else confirm though?

OpenStudy (princessaurora):

Thanks a lot, Is there a more simplified version of the answer or is that as far as I can simplify it?

OpenStudy (princessaurora):

like is that the correct answer form?

OpenStudy (johnweldon1993):

Well no they want the answer rounded to the nearest hundredth! So we actually must evaluate that \[\large SE = 0.57\]

OpenStudy (princessaurora):

Yeah I got that in my calculator just double checking thanks a lot.

OpenStudy (johnweldon1993):

Of course! :)